NI ring alignment shift 28 Jan 2009

Go down

NI ring alignment shift 28 Jan 2009

Post  Alessio on Thu Jan 29, 2009 3:23 pm

logbook entry evening 28Jan2009

PR misaligned by -1000

Baseline used 917199035

NI Central spot switched on

-NI BR moved 500 forward to 4525 Time: Jan28,09-18:53:07 > GPS: 917200402
-NI BR moved 500 forward to 5025 forward Time: Jan28,09-18:54:56 > GPS: 917200511
-NI BR moved 500 backward to 4525 Time: Jan28,09-18:57:45 > GPS: 917200680

-rotor moved -96 to -76 18h01
now we have 140mW for central spot

-NI BR moved 500 backward to 4025 Time: Jan28,09-19:04:33 > GPS: 917201088

Ni Central spot switch off Time: Jan28,09-19:29:06 > GPS: 917202560

This test allows us to deduce the directions for horizontal motion:
NI BR motor POSITIVE > Movement in phase camera NEGATIVE > B7p movement NEGATIVE
Movement in phase camera of 9 pixels > Movement of NI UL motor 500 steps

From experiment done during previous shift (21850 > Fig 5 ) we can deduce the directions for Vertical motion:
NI UL motor POSITIVE > Movement in phase camera NEGATIVE > B7p movement POSITVE
Movement in phase camera of 20 pixels > Movement of NI UL motor 1000 steps

NOTE: this test was done using the central spot (positive lens). If we are using the ring (negative lens) then B7p movement is the opposite.

If we consider the TCS mirror calibration of 67steps/mm then we can estimate the correspondance between position on the input mirror and position in the phase camera image to be 0.75mm/pixel.
We can double check this by using the measurement of the waist of channel V1:PC1_LSB_waistx and V1:PC1_LSB_waisty which is about 20 pixels. As we know that the waist of the beam
at the input mirror is about 20mm then we can make a second estimation of 1mm/pixel.

Based on measured positions of central spot in phase image relative to LSB amplitude position corrective movements of mirror made:
NI UL moved back 200 to 501
NI BR moved forward 75 to 4100

switch on ring to heat mirror and use drift of B7p to determine if centred
Power set to 4.5 W
shutter switched on Time: Jan28,09-19:52:18 > GPS: 917203953
shutter switched off Time: Jan28,09-19:57:17 > GPS: 917204252

Power set to 12 W
shutter switched on Time: Jan28,09-20:06:26 > GPS: 917204801
shutter switched off Time 19h10-UTC

The B7p positions moves very little suggesting that the ring is centred on the mirror.
However the phase camera image continues to indicate mis-centering.This is not well understood.
Motor positions are NI-UL = 501, NI-BR = 4100, NI-BL = 200

19h21-UTC realign PR due to thermal effects making B7p drift.

New baseline taken at 917211453 as a tilt seems to have been introduced in the phase image

Power set to 12 W
shutter switched on Time: Jan28,09-22:01:58 > GPS: 917211733
shutter switched off Time: Jan28,09-22:05:52 > GPS: 917211967

B7p_posx was seen to rise slightly so we adjust the motor to compensate.
NI BR moved backward 500 to 3600 Time: Jan28,09-22:23:40 > GPS: 917213035

We used the B7p position to center the CO2 ring.
The centred ring (using this technique) is almost in the same place as the centred central spot.
The phase camera image makes sense with the central spot but is not at all understood when using the ring. More simulation is perhaps needed to better understand.


These measurements confirm our suggestions that the centre of the ring should correspond to the position of the central spot, given that the two beams are well aligned. This was also suggested in the old measurements when the bright spot of the ring was clearly visible in the phase camera images almost exactly in the same position as the central spot.

Posts : 127
Join date : 2008-11-07
Age : 46
Location : Rome

View user profile

Back to top Go down

Back to top

- Similar topics

Permissions in this forum:
You cannot reply to topics in this forum